# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix D Proof of Lipschitz-continuity Bound
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\bigg{[}r_{\Delta}(s_{g}^{\prime},s_{\Delta}^{\prime},\widetilde{a}_{g},\widetilde{a}_{\Delta})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},a_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})\bigg{]}$
$-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\bigg{[}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{g},\widetilde{a}_{\Delta^{\prime}})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},s{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
@@
$\leq\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}r_{\Delta}(s_{g}^{\prime},s_{\Delta}^{\prime},\widetilde{a}_{g},\widetilde{a}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{g},\widetilde{a}_{\Delta^{\prime}})\bigg{|}$
@@
$+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})$
@@
$-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})\bigg{|}$
@@
$+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim{\widetilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}^{\prime})\sim{\widetilde{\mathcal{J}}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
@@
The first equality rewrites the equations with their respective maximizing actions. The first inequality upper-bounds this difference by allowing all terms to share the common action $\widetilde{a}$. Using the Bellman equation, the second equality expands $\widehat{Q}_{k}^{T+1}$ and $\widehat{Q}_{k^{\prime}}^{T+1}$. The second inequality follows from the triangle inequality. The third equality follows by subtracting the common $r_{g}(s_{g}^{\prime},\widetilde{a_{g}})$ terms from the reward and noting that the two expectation terms $\mathbb{E}_{s_{g}^{\prime},s_{\Delta}^{\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},\widetilde{a}_{g},s_{\Delta},\widetilde{a}_{\Delta})}$ can be combined into a single expectation $\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime}\sim\widetilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}$.
@@
In the special case where $a_{\Delta}=\widetilde{a}_{\Delta}$, we derive a closed form expression for $\widetilde{\mathcal{J}}_{k}$ in Lemma D.10. To justify the third inequality, the second term follows from the induction hypothesis and the first term follows from the following derivation:
@@
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})\bigg{|}$
@@
$\quad\quad\quad=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}r^{\widetilde{a}}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}^{\widetilde{a}}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
@@
$\quad\quad\quad\leq\|r_{l}^{\widetilde{a}}(\cdot,\cdot)\|_{\infty}\cdot\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathrm{TV}(\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}})$
$\quad\quad\quad\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
@@
The above derivation follows from the same argument as in the base case where $r_{l}^{\widetilde{a}}(s_{i}^{\prime},s_{g}^{\prime}):=r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})$ for any $i\in\Delta\cup\Delta^{\prime}$. Similarly, if (I)$<$(II), an analogous argument that replaces $\widetilde{a}_{\Delta}$ with $\widehat{a}_{\Delta}$ yields the same result.
@@
Remark D.6.
Given a joint transition probability function $\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}$ as defined in Definition D.4, we can recover the transition function for a single agent $i\in\Delta\cup\Delta^{\prime}$ given by $\mathcal{J}_{1}$ using the law of total probability and the conditional independence between $s_{i}$ and $s_{g}\cup s_{[n]\setminus i}$ in Eq. 29. This characterization is crucial in Lemma D.7 and Lemma D.5:
@@
Here, we use a conditional independence property proven in Lemma D.9.
Lemma D.7.
$\mathrm{TV}\bigg{(}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}\bigg{)}\leq\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$

### v3

@@ Appendix E Proof of Lipschitz-continuity Bound
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\tilde{a}_{g},s_{\Delta}^{\prime},\tilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\hat{Q}_{k}^{T}(s^{\prime\prime}_{g},a_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})\bigg{]}$
@@
$\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\bigg{[}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\tilde{a}_{g},\tilde{a}_{\Delta^{\prime}})$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\tilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\tilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\hat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},s{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
$\leq\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}r_{\Delta}(s_{g}^{\prime},s_{\Delta}^{\prime},\tilde{a}_{g},\tilde{a}_{\Delta})$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\tilde{a}_{g},\tilde{a}_{\Delta^{\prime}})\bigg{|}$
@@
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\tilde{a}_{g},s_{\Delta}^{\prime},\tilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\hat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})$
$\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\tilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\tilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\hat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})\bigg{|}$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim{\tilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\hat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}^{\prime})\sim{\tilde{\mathcal{J}}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\hat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
@@
The first equality rewrites the equations with their respective maximizing actions. The first inequality upper-bounds this difference by allowing all terms to share the common action $\tilde{a}$. Using the Bellman equation, the second equality expands $\hat{Q}_{k}^{T+1}$ and $\hat{Q}_{k^{\prime}}^{T+1}$. The second inequality follows from the triangle inequality. The third equality follows by subtracting the common $r_{g}(s_{g}^{\prime},\tilde{a_{g}})$ terms from the reward and noting that the two expectation terms $\mathbb{E}_{s_{g}^{\prime},s_{\Delta}^{\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},\tilde{a}_{g},s_{\Delta},\tilde{a}_{\Delta})}$ can be combined into a single expectation $\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime}\sim\tilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}$.
@@
In the special case where $a_{\Delta}=\tilde{a}_{\Delta}$, we derive a closed form expression for $\tilde{\mathcal{J}}_{k}$ in Lemma E.11. To justify the third inequality, the second term follows from the induction hypothesis and the first term follows from the following derivation:
@@
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})}{k}-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})}{k^{\prime}}\bigg{|}$
@@
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{\sum_{i\in\Delta}r^{\tilde{a}}_{l}(s_{i}^{\prime},s_{g}^{\prime})}{k}-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{\sum_{i\in\Delta^{\prime}}r_{l}^{\tilde{a}}(s_{i}^{\prime},s_{g}^{\prime})}{k^{\prime}}\bigg{|}$
@@
$\leq\|r_{l}^{\tilde{a}}\|_{\infty}\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathrm{TV}(\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}})$
@@
The above derivation follows from the same argument as in the base case where $r_{l}^{\tilde{a}}(s_{i}^{\prime},s_{g}^{\prime}):=r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})$ for any $i\in\Delta\cup\Delta^{\prime}$. Similarly, if (I) $<$ (II), an analogous argument that replaces $\tilde{a}_{\Delta}$ with $\hat{a}_{\Delta}$ yields the same result.
@@
Remark E.6.
Given a joint transition probability function $\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}$ as defined in Definition E.4, we can recover the transition function for a single agent $i\in\Delta\cup\Delta^{\prime}$ given by $\mathcal{J}_{1}$ using the law of total probability and the conditional independence between $s_{i}$ and $s_{g}\cup s_{[n]\setminus i}$ in Equation 26. This characterization is crucial in Lemma E.7 and Lemma E.5:
@@
Here, we use a conditional independence property proven in Lemma E.10.
Lemma E.7.
The TV-distance between the next-step expected empirical distribution functions is bounded by the TV-distance between the existing empirical distribution functions.
$\mathrm{TV}\bigg{(}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}\bigg{)}$
$\quad\quad\quad\quad\quad\quad\leq\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$